The sides of and equilateral triangle are 8 inches each. What is the height of the triangle?
Answer by checkley71(8403)

HERE YOU HAVE A RIGHT TRIANGLE WITH BASE=4 & HYPOTENUSE=8 & THE HEIGHT (X) BEING THE UNKNOWN SIDE.
4^2+X^2=8^2
16+X^2=64
X^2=64-16
X^2=48
X=SQRT48
X=6.93 ANSWER FOR THE HEIGHT.